# Data Science Principles for Interpretable and Explainable AI

### v1

@@ Abstract
Society’s capacity for algorithmic problem-solving has never been greater. Artificial Intelligence is now applied across more domains than ever, a consequence of powerful abstractions, abundant data, and accessible software. As capabilities have expanded, so have risks, with models often deployed without fully understanding their potential impacts. Interpretable and interactive machine learning aims to make complex models more transparent and controllable, enhancing user agency. This review synthesizes key principles from the growing literature in this field.
We first introduce precise vocabulary for discussing interpretability, like the distinction between glass box and explainable algorithms. We then explore connections to classical statistical and design principles, like parsimony and the gulfs of interaction. Basic explainability techniques – including learned embeddings, integrated gradients, and concept bottlenecks – are illustrated with a simple case study. We also review criteria for objectively evaluating interpretability approaches. Throughout, we underscore the importance of considering audience goals when designing interactive algorithmic systems. Finally, we outline open challenges and discuss the potential role of data science in addressing them. Code to reproduce all examples can be found at https://go.wisc.edu/3k1ewe.
@@ 1 Introduction
The success of Artificial Intelligence (AI) stems in part from its ability to abstract away details of its context. When an algorithm is applied to predict responses from a collection of features, it matters little whether the response is the shape of a galaxy or the next word in a sentence. Such abstraction has enabled the design of versatile methods with remarkably diverse applications. Nonetheless, algorithms are never used in a vacuum. Models affect human well-being, and it is important to systematically study the human-algorithm interface. How can we ensure that an algorithm behaves acceptably in unforeseen circumstances? How can we prevent harm and unfair treatment111As Recht (2023) writes only half-jokingly, “All models are wrong, but some are dangerous.”? How can align algorithmic behavior with our expectations? The field of interpretable machine learning and explainable AI (XAI) have emerged to address these questions, helping develop techniques that allow algorithmic abstractions to be used wisely within messy, real-world contexts. This literature has engaged researchers from various backgrounds, from theorists seeking to characterize the fundamental limits of explainability techniques (Williamson and Feng, 2020; Bilodeau et al., 2024) to philosophers and psychologists critiquing popular discourses surrounding them (Krishnan, 2019; Nussberger et al., 2022).
@@
Caruana et al. (2015) aimed to help a hospital triage incoming pneumonia patients, ensuring that high-risk patients are admitted while allowing lower-risk patients to receive outpatient care. Using a dataset of past pneumonia patients, their model achieved an AUC of 0.857 in predicting patient mortality based on lab and medical profiles. Surprisingly, an important feature analysis revealed that asthmatic patients had lower predicted probability of death. This arose because the original training data were observational. Historically, asthmatic patients were treated more aggressively than those without asthma, and therefore enjoyed lower mortality. Directly deploying such a model would put asthma patients at risk. Fortunately, their model was editable, so the asthma association could be removed before deployment.
@@
Gu et al. (2019) designed stickers that, when attached to stop signs, caused object detector algorithms used by self-driving cars to misclassify them as speed limits. This real-world adversarial attack exploits the sensitivity of algorithms to small (and, to humans, seemingly arbitrary) input perturbations. More interpretable models could yield systems with more predictable behavior.
@@
In Google’s first public demo of the generative AI chatbot Bard, it was asked, “What new discoveries from the James Webb Space Telescope (JWST) can I tell my 9 year old about?” The chatbot responded with three seemingly legitimate facts. Only after the demo was shared online did astronomers point out that the third fact, “JWST took the very first pictures of a planet outside of our own solar system,” was in fact untrue — the first exoplanet had been imaged in 2004 (Kundaliya, 2023). AI chatbots can “hallucinate” falsehoods, even when their output seems authoritative. Research on XAI could shed light on how these systems generate responses, allowing them to be treated less like an oracle and more like a search engine.
@@
These examples beg the question: What makes an algorithm interpretable? Directly answering this questions is complex, so let’s consider a simpler one: What makes a data visualization effective? This question has been studied for decades (Sedlmair et al., 2012; Agrawala et al., 2011; Tufte, 2001; Cleveland, 1993), and every data scientist has firsthand experience with it. A good visualization streamline a taxing cognitive operation into a perceptual one. It helps when the visual encoding – the mapping from sample properties to graphical elements – uses representations that are already familiar or easily learnable. Further, the graphical elements must be legible and well-annotated. We also tend to learn more information-dense visualizations (Tufte, 2001; Oppermann and Munzner, 2022), since they prevent oversummarization and can highlight details for follow-up study. Similarly, an interpretable model can be broken down into relevant components, each of which can be assigned meaning. Instead of information density (showing more of the data), interpretability relies on faithfulness (showing more of the model).
@@
The parallels run deeper. As in a good visualization, the data provenance of a trustworthy model can be traced back to the original measurement mechanism — beautiful design and high-accuracy have little value otherwise. Moreover, like visualization, interpretability must be tailored to its audience and the tasks it is designed to solve (Lipton, 2018). There are different levels of data literacy, and visual representations may be familiar to some audiences but not others. Similarly, AI models are employed across a range of problem domains, necessitating validation in realistic settings (see Section 3). Finally, effective visualizations push readers beyond passive consumption — they inspire deeper exploration of complexity. Likewise, interpretability can support the transition from automated to augmented decision-making (Heer, 2019), enhancing rather than substituting human reason.
@@
Algorithmic interpretability can be approached with the same nuance that is already routine in data visualization. The data visualization literature has identified systematic design approaches and formal evaluation criteria, cautioning against blanket statements about entire approaches222Even information density has been thoughtfully critiqued (Borkin et al., 2013).. Similarly, discussions of interpretability can go beyond dichotomies about models being either glass or black boxes. To help us navigate the gray areas of this growing field, Section 1.1 introduces relevant vocabulary. Section 2 outlines representative techniques and applies them to a simple simulation example. This section, and the code that accompanies it (https://go.wisc.edu/3k1ewe), can serve as a tutorial on the practical implementation of interpretability techniques. Finally, Sections 3–4 examine the conceptual and technical questions required for effective evaluation and progress as the AI landscape evolves.
@@ 1.1 Vocabulary
The terms used in interpretability research can be confusing, because though terms like “explainable” are familiar in a colloquial sense, they have a precise technical meaning in the literature. Indeed, one difficulty is that many people agree that algorithmic interpretability is important — they just have different ideas for what it might look like. Therefore, it helps to have vocabulary that distinguishes key properties while maintaining an appropriate level of abstraction.
@@
An important distinction is between an interpretable model and an explainability technique (Murdoch et al., 2019; Rudin, 2019). An interpretable model is one that, by virtue of its design, is easy to accurately describe and alter. For this reason, they are often called glass boxes. A canonical example is a sparse linear model. For any new example, a prediction can be constructed in a single pass over the features with nonzero coefficients. In contrast, explainability techniques are designed to improve our mental models of arbitrary black box models. A common example is the partial dependence plot (Friedman, 2001). Though originally designed to summarize gradient boosting fits , the fact that it only requires pairs means it applies to any supervised learner. If an interpretable model is a glass box whose inner workings are transparent, then an explainability technique is a systematic way of analyzing the outputs emerging from a black box.
@@
We can relate these approaches to data sciences and design principles; see also Table 1. First, consider interpretable models. When we call a linear model interpretable, we are invoking parsimony and simulatability. Parsimony means that predictions can be traced back to a few model components, each of which comes with a simple story attached. For example, sparse linear models have few nonzero coefficients, and the relationship between each coefficient and the output can be concisely described. A similar principle applies to generalized additive modeling, which have few allowable interactions (Caruana et al., 2015), or in latent variable models, which have few underlying factors (Sankaran and Holmes, 2023). Simulatability formalizes the idea that predictions can be manually reconstructed from model descriptions. For example, in decision trees and falling rules lists, this can be done by answering a sequence of yes-no questions. Finally, since interpretation requires human interaction, it can be affected by the gulfs of interaction (Hutchins et al., 1985). Ideally, an interpretability method should allow users to quickly query properties of the underlying model. Further, the method’s outputs should be immediately relevant to the query, requiring no further cognitive processing. To the extent that a method has reached these ideals, it has reduced the gulfs of execution and evaluation.
@@
Unfortunately, even when restricting focus to these glass box models, these definitions are still somewhat ambiguous. These gaps lead to larger questions about evaluation. For example, simulatability alone does not quantify the user effort required to form predictions; it also ignores their accuracy. Linear models become unwieldy when they have many nonzero coefficients, and deep decision trees take time to parse. Similarly, latent variables are only interpretable if they can be related to existing knowledge.
@@
Among explainability techniques, the key question is scale, and this leads to the distinction between global and local explanations. Global explanations summarize the overall structure of a model. Partial dependence profiles are an example of a global explanation, because they describe across the range of possible inputs. In contrast, local explanations support reasoning about individual predictions. For example, when a sentiment analysis model classifies a particular hotel review as negative, we might be curious about which phrases were most important in that classification. Global explanations shed light on the high-level properties of a model while local explanations dissect specific instances of algorithmic judgment. Local explanations are built on the observation that perturbations, either of the model or the data, provide valuable context for understanding algorithmic decisions. Sensitivity, a principle with a long history in statistics (Tukey, 1959; Huber, 1964; Holmes, 2017), is a fruitful path to explainability. Nonetheless, as above, we caution that differences between local and global can blur in practice. For example, concept bottleneck models (Section 2.3.4), globally restricts the space of possible prediction functions, forcing them to pass through an interpretable concept bottleneck. However, editing the bottleneck for specific examples and support local interventions, ensuring that concepts like race or gender are not used downstream. In light of these ambiguities, it is helpful to think of this vocabulary as highlighting the main axes of variation in the space of techniques, a checklist that can guide application and evaluation of interpretability in practice.
@@
Simulatability | Given a sample and a model description, how easy would it be for a user to manually derive the associated prediction? Regression models with few coefficients and algorithms like rule lists tend to be more simulatable than those that involve more intensive or multistep pooling of evidence. |
Sensitivity | How robust are predictions to small changes in either the data or the model? Ideas from robust statistics can inform how we understand variable importance and stability in more complex AI algorithms. |
@@ 2 Methods
How is interpretability put into practice? We next review approaches from the direct interpretability and XAI perspectives. Rather than attempting to review all proposals from this rapidly growing literature, we prioritize the simple but timeless ideas that lie behind larger classes of methods. To ground the discussion, we include a simulation example where the generative mechanism is simple to communicate, but complex enough to warrant analysis with AI.
@@ 2.1 Simulation design
The simulation is inspired by the types of longitudinal studies often encountered in microbiome data science. Consider following 500 participants for two months, gathering microbiome samples each day. At the study’s conclusion, each participant has their health checked, and they are assigned to either healthy or disease groups. For example, the participants may have contracted HIV or developed Type I Diabetes, as in (Gosmann et al., 2017) and (Kostic et al., 2015), respectively. The question of interest is whether there are systematic differences in the microbiome trajectories between groups. These might have diagnostic potential and could shed light on the underlying mechanisms of disease development. For example, does the disease group have lower levels of a protective species? Do they exhibit unstable dynamics for any important microbial subcommunities? For either fundamental biology or diagnostic potential, any models applied to the data be either interpretable or explainable – classification alone has little value unless it can be incorporated into a medically coherent narrative.
In our hypothetical scenario, we model participants across time points, with each sample containing abundances of microbial species. Figure 1 illustrates trajectories for a selection of taxa and participants. The generative mechanism is as follows: we create community trajectories , for , by following these steps:
@@
To generate class assignments , we assign the community weights to 24 clusters via means. These clusters are then randomly divided into 12 healthy and 12 disease groups. This approach therefore links subjects with similar underlying community trajectories into similar health outcomes in a way that is highly nonlinear.
In practice, the steps of the generative mechanism would not be known. Visually examining example trajectories in Figure 1d does not suggest clear choices about which microbiome features are relevant for disease classification. Each taxon exhibits a complex, nonlinear trend over time, and while some participants show “blooms” at similar times, it’s unclear whether these are linked to disease. Further, the sheer number of taxa makes drawing conclusions through manual inspection impractical. Algorithmic approaches are necessary to distinguish disease from healthy trajectories and identify taxonomic or temporal features of interest, which could then be validated in follow-up studies involving direct interventions on features in a controlled setting.
2.2 Directly interpretable models
We can often learn salient, non-obvious characteristics of the data by applying a directly interpretable model. Some have argued that with sufficient ingenuity, such models can rival the accuracy and efficiency of any black box (Rudin, 2019). Even otherwise, the results of directly interpretable analyses can guide the application of more sophisticated models later on. With this in mind, we briefly review sparse logistic regression (Friedman et al., 2010) and decision trees (Loh, 2014), apply them to the simulation, and interpret the results.

### v2

@@ Abstract
Society’s capacity for algorithmic problem-solving has never been greater. Artificial Intelligence is now applied across more domains than ever, a consequence of powerful abstractions, abundant data, and accessible software. As capabilities have expanded, so have risks, with models often deployed without fully understanding their potential impacts. Interpretable and interactive machine learning aims to make complex models more transparent and controllable, enhancing user agency. This review synthesizes key principles from the growing literature in this field. We first introduce precise vocabulary for discussing interpretability, like the distinction between glass box and explainable models. We then explore connections to classical statistical and design principles, like parsimony and the gulfs of interaction. Basic explainability techniques – including learned embeddings, integrated gradients, and concept bottlenecks – are illustrated with a simple case study. We also review criteria for objectively evaluating interpretability approaches. Throughout, we underscore the importance of considering audience goals when designing interactive data-driven systems. Finally, we outline open challenges and discuss the potential role of data science in addressing them. Code to reproduce all examples can be found at https://go.wisc.edu/3k1ewe.
doi:
0000keywords:
Research Article [type=corresp,id=c1]Corresponding author.
@@ 1 Introduction
The success of Artificial Intelligence (AI) stems in part from its ability to abstract away details of its context. When a model is applied to predict responses from a collection of features, it matters little whether the response is the shape of a galaxy or the next word in a sentence. Such abstraction has enabled the design of versatile methods with remarkably diverse applications. Nonetheless, models are never used in a vacuum. Models affect human well-being, and it is important to systematically study the human-model interface. How can we ensure that a model behaves acceptably in unforeseen circumstances? How can we prevent harm and unfair treatment? How can we align model behavior with our expectations? The field of interpretable machine learning and explainable AI (XAI) has emerged to address these questions, allowing model-based abstractions to be used wisely within messy, real-world contexts. This literature has engaged researchers from various backgrounds, from theorists seeking to characterize the fundamental limits of explainability techniques (Williamson and Feng, 2020; Bilodeau et al., 2024) to philosophers and psychologists critiquing popular discourses surrounding them (Krishnan, 2019; Nussberger et al., 2022).
@@
Caruana et al. (2015) aimed to help a hospital triage incoming pneumonia patients, ensuring that high-risk patients are admitted while allowing lower-risk patients to receive outpatient care. Using a dataset of past pneumonia patients, their model achieved an area under the receiver operating characteristic curve (AUC) of 0.857 in predicting patient mortality based on lab and medical profiles. Surprisingly, an important feature analysis revealed that asthmatic patients had lower predicted probability of death. This arose because the original training data were observational. Historically, asthmatic patients were treated more aggressively than those without asthma, and therefore enjoyed lower mortality. Directly deploying such a model would put asthma patients at risk. Fortunately, their model was editable, so the asthma association could be removed before deployment.
@@
Gu et al. (2019) designed stickers that, when attached to stop signs, caused object detector models used by self-driving cars to misclassify them as speed limits. This real-world adversarial attack exploits the sensitivity of models to small (and, to humans, seemingly arbitrary) input perturbations. More interpretable models could yield systems with more predictable behavior.
@@
In Google’s first public demo of the generative AI chatbot Bard, it was asked, “What new discoveries from the James Webb Space Telescope (JWST) can I tell my 9 year old about?” The chatbot responded with three seemingly legitimate facts. Only after the demo was shared online did astronomers point out that the third fact, “JWST took the very first pictures of a planet outside of our own solar system,” was in fact untrue — the first exoplanet had been imaged in 2004 (Kundaliya, 2023). AI chatbots can “hallucinate” falsehoods, even when their output seems authoritative. Research on XAI could shed light on how these systems generate responses, allowing them to be treated less like oracles and more like search engines.
@@
These examples beg the question: What makes a model interpretable? Directly answering this questions is complex, so let’s consider a simpler one: What makes a data visualization effective? This question has been studied for decades (Cleveland, 1993; Tufte, 2001; Agrawala et al., 2011; Sedlmair et al., 2012), and every data scientist has firsthand experience with it. A good visualization streamlines a taxing cognitive operation into a perceptual one. It helps when the visual encoding – the mapping from sample properties to graphical elements – uses representations that are already familiar or easily learnable. Further, the graphical elements must be legible and well-annotated. We also tend to learn more from information-dense visualizations (Tufte, 2001; Oppermann and Munzner, 2022), since they prevent oversummarization and can highlight details for follow-up study. Similarly, an interpretable model can be broken down into relevant components, each of which can be assigned meaning. Instead of information density (showing more of the data), interpretability relies on faithfulness (showing more of the model).
@@
The parallels run deeper. As in a good visualization, the data provenance of a trustworthy model can be traced back to the original measurement mechanism — beautiful design and high-accuracy have little value otherwise. Moreover, like visualization, interpretability must be tailored to an audience and the problems they need solved (Lipton, 2018). There are different levels of data literacy, and visual representations may be familiar to some audiences but not others. Similarly, AI models are employed across a range of problem domains, necessitating validation in realistic settings (see Section 3). Finally, effective visualizations push readers beyond passive consumption — they inspire deeper exploration of complexity. Likewise, interpretability can support the transition from automated to augmented decision-making (Heer, 2019), enhancing rather than substituting human reason.
@@
Model interpretability can be approached with the same nuance that is already routine in data visualization. The data visualization literature has identified systematic design approaches and formal evaluation criteria, cautioning against blanket statements about entire approaches. Indeed, even information density has been thoughtfully critiqued (Borkin et al., 2013). Similarly, discussions of interpretability can go beyond dichotomies about models being either glass or black boxes. To help us navigate the gray areas of this growing field, Section 1.1 introduces relevant vocabulary. Section 2 outlines representative techniques and applies them to a simple simulation example. This section, and the code that accompanies it (https://go.wisc.edu/3k1ewe), can serve as a tutorial on the practical implementation of interpretability techniques. Finally, Sections 3 – 4 examine the conceptual and technical questions required for effective evaluation and progress as the AI landscape evolves.
@@ 1.1 Vocabulary
The terms used in interpretability research can be confusing, because though terms like “explainable” are familiar in a colloquial sense, they have a precise technical meaning in the literature. Indeed, one difficulty is that many people agree that model interpretability is important — they just have different ideas about what it looks like. Therefore, it helps to have vocabulary that distinguishes key properties while maintaining an appropriate level of abstraction.
@@
An important distinction is between an intrinsically interpretable model and an explainability technique (Murdoch et al., 2019; Rudin, 2019). An intrinsically interpretable model is one that, by virtue of its design, is easy to accurately describe and alter. For this reason, they are often called glass boxes. A canonical example is a sparse linear model. For any new example, a prediction can be constructed in a single pass over the features with nonzero coefficients. In contrast, explainability techniques are designed to improve our mental models of arbitrary black box models. A common example is the partial dependence profile (Friedman, 2001), defined as , which describes the influence of the feature when holding all others fixed at values observed in the original data . Though originally designed to summarize gradient boosting fits , the fact that it only requires pairs means it applies to any supervised learner. To summarize, interpretable models are like glass boxes whose inner workings are transparent, while explainability techniques are systematic ways of analyzing the outputs emerging from black boxes.
@@
We can relate these approaches to data science and design principles; see Table 1. First, consider intrinsically interpretable models. When we call a linear model interpretable, we are invoking parsimony and simulatability. Parsimony means that predictions can be traced back to a few model components, each of which comes with a simple story attached. For example, sparse linear models have few nonzero coefficients, and the relationship between each coefficient and the output can be concisely described. A similar principle applies to generalized additive modeling, which have few allowable interactions (Caruana et al., 2015), or in latent variable models, which have few underlying factors (Sankaran and Holmes, 2023). Simulatability formalizes the idea that predictions can be manually reconstructed from model descriptions. For example, in decision trees and falling rules lists, this can be done by answering a sequence of yes-no questions. Finally, since interpretation requires human interaction, it can be affected by the gulfs of interaction (Hutchins et al., 1985). Ideally, an interpretability method should allow users to quickly query properties of the underlying model. Further, the method’s outputs should be immediately relevant to the query, requiring no further cognitive processing. To the extent that a method has reached these two ideals, it has reduced the gulfs of execution and evaluation, respectively.
@@
Unfortunately, even when restricting focus to glass box models, these definitions are still somewhat ambiguous, leading to issues with evaluation. For example, simulatability alone does not quantify the user effort required to form predictions; it also ignores their accuracy. Linear models become unwieldy when they have many nonzero coefficients (Poursabzi-Sangdeh et al., 2021), and deep decision trees take time to parse. Similarly, latent variables are only interpretable if they can be related to existing knowledge.
@@
Among explainability techniques, a key question is scale, and this leads to the distinction between global and local explanations (Lundberg et al., 2020). Global explanations summarize the overall structure of a model. Partial dependence profiles are an example of a global explanation, because they describe across the range of possible inputs. In contrast, local explanations support reasoning about an individual sample’s prediction. For example, when a sentiment analysis model classifies a particular hotel review as negative, we might be curious about which phrases were most important in that classification. Global explanations shed light on the high-level properties of a model while local explanations dissect specific instances of a model’s prediction. Local explanations are built on the observation that perturbations, either of the model or the data, provide valuable context for understanding model-based decisions. Sensitivity, a principle with a long history in statistics (Tukey, 1959; Huber, 1964; Holmes, 2017), is a fruitful path to explainability. Nonetheless, we caution that differences between local and global can blur in practice (Dandl et al., 2023; Herbinger et al., 2022). For example, concept bottleneck models (Section 2.3.4), globally restrict the space of possible prediction functions, forcing them to pass through an interpretable concept bottleneck. However, editing the bottleneck for specific examples allows for local interventions. As detailed in Section 2.3.4, if sensitive attributes are used as concepts, then their coefficients can be set to zero, ensuring that they don’t influence the final prediction, even indirectly through related features. In light of these ambiguities, it is helpful to think of the vocabulary in this section as highlighting the main axes of variation in the space of techniques, a map to guide application and evaluation of interpretability methods.
@@
Simulatability | Given a sample and a model description, how easy would it be for a user to manually derive the associated prediction? Regression models with few coefficients and approaches like rule lists tend to be more simulatable than those that involve more intensive or multistep pooling of evidence. |
Sensitivity | How robust are predictions to small changes in either the data or the model? Ideas from robust statistics can inform how we understand variable importance and stability in more complex AI models. |
@@ 2 Methods
How is interpretability put into practice? We next review approaches for direct interpretability and XAI. Rather than attempting to review all proposals from this rapidly growing literature, we prioritize the simple but timeless ideas that lie behind larger classes of methods. To ground the discussion, we include a simulation example where the generative mechanism is simple to communicate, but complex enough to warrant analysis with AI.
@@ 2.1 Simulation design
The simulation is inspired by the types of longitudinal studies often encountered in microbiome data science (Jeganathan et al., 2018; Silverman et al., 2018; Kodikara et al., 2022). Consider following 500 participants for two months, gathering microbiome samples each day. At the study’s conclusion, each participant has their health checked, and they are assigned to either healthy or disease groups. For example, the participants may have contracted HIV or developed Type I Diabetes, as in (Gosmann et al., 2017) and (Kostic et al., 2015), respectively. The question of interest is whether there are systematic differences in the microbiome trajectories between groups. These might have diagnostic potential and could shed light on the underlying mechanisms of disease development. For example, does the disease group have lower levels of a protective species? Do they exhibit unstable dynamics for any important microbial subcommunities? For either fundamental biology or diagnostic potential, any models applied to the data must be either interpretable or explainable – classification alone has little value unless it can be incorporated into a medically coherent narrative.
In our hypothetical scenario, we simulate participants across time points, with each sample containing abundances of microbial species. Figure 1 illustrates trajectories for a selection of taxa and participants. The generative mechanism is as follows: we create community trajectories , for by sampling:
@@
To generate class assignments , we cluster the community weights into 24 clusters. These clusters are then randomly divided into 12 healthy and 12 disease groups. Note that each cluster contains a mixture of the community trajectories. This approach therefore links subjects with similar underlying community trajectories into similar health outcomes in a way that is highly nonlinear.
In practice, the steps of the generative mechanism would not be known. Visually examining example trajectories in Figure 1d does not suggest clear choices about which microbiome features are relevant for disease classification. Each taxon exhibits a complex, nonlinear trend over time, and while some participants show “blooms” at similar times, it’s unclear whether these are linked to disease. Further, the sheer number of taxa makes drawing conclusions through manual inspection impractical. Computational approaches are necessary to distinguish disease from healthy trajectories and identify taxonomic or temporal features of interest, which could then be validated in follow-up studies involving direct interventions on features in a controlled setting.
2.2 Intrinsically interpretable models
We can often learn salient, non-obvious characteristics of the data by applying a intrinsically interpretable model. Some have argued that with sufficient ingenuity, such models can rival the accuracy and efficiency of any black box (Rudin, 2019). Even otherwise, the results of intrinsically interpretable analyses can guide the application of more sophisticated models later on. With this in mind, we briefly review sparse logistic regression (Friedman et al., 2010) and decision trees (Loh, 2014), apply them to the simulation, and interpret the results.